OpenStudy (anonymous):

please help me with algebra 2 work??

OpenStudy (studygurl14):

Use the quadratic thing, or what it is called: \[x = \frac{ -b \pm \sqrt{b^{2} - 4ac} }{ 2a }\]

OpenStudy (anonymous):

can you take me through it step by step?

OpenStudy (studygurl14):

\[ax^{2} + bx + c = 0\] \[-x^{2} + 5x = 3\] \[-x^{2} + 5x -3 = 0\] So.... a = -1 b = 5 and c = -3 Plug in values of a, b, and c into the quadratic formula then solve

OpenStudy (studygurl14):

\[x = \frac{ -b \pm \sqrt{b^{2} - 4ac} }{ 2a }\] \[x = \frac{ -(5) \pm\sqrt{(5)^{2} - 4(-1)(-3)} }{ 2(-1) }\]

OpenStudy (anonymous):

so which one would be?

OpenStudy (studygurl14):

you should be able to do the rest to find the answer. I'm not going to solve it all for you. I'll help only if you are confused about what to do or how to do something. Besides, I have to leave soon. :)